# The Amaterasu particle: constraining the superheavy dark matter origin of UHECRs

### v1

@@ III.1 Present UHE telescopes
In this sub-section, we present the constraints on SHDM lifetime from different observations of UHECRs, UHE gamma-rays and UHE neutrinos and analyse the origin of the TA Amaterasu event. The constraints are obtained by normalising the flux of these particles with the experimental data or limits from non-observation. For UHECRs, we consider the spectrum measured by TA together with the recent TA Amaterasu event as depicted in the left plot of Fig. 1. Including the UHECR data from other experiments such as PAO, AGASA and HiRes will yield similar values of DM lifetime as the CR spectral shapes from these experiments are also very similar to that of TA. For the UHE gamma-rays, we take the upper limits on the flux from TA [53, 54] and PAO [84, 49]. Whereas for the UHE neutrinos, we consider the upper limits on the flux from PAO [98] and IceCube [58]. We consider a wide range of $m_{\chi}=(10^{11}-10^{15})$ GeV and different two-body decay modes. For the decay modes, we take all the quarks (u, d, s, b, t, c), the gauge bosons (W, Z, $\gamma$, gluon), and the Higgs boson (h). The lower limit of $m_{\chi}$ is chosen to be $10^{11}$ GeV as our primary goal is to analyse the origin of the TA Amaterasu event at $2.44\times 10^{11}$ GeV. On the other hand, the upper limit of $m_{\chi}=10^{15}$ GeV is taken to be close to the GUT scale. For the given $m_{\chi}$, we constrain the lifetime such that the UHE secondary particle observations, limits and sensitivities are satisfied. In particular, we follow convention of exclusion plots and constrain the inverse of the lifetime, i.e., $\tau_{\chi}^{-1}$ rather than the lifetime ($\tau_{\chi}$). The constraints on $\tau_{\chi}^{-1}$ as a function of $m_{\chi}$ for different decay modes are depicted in Fig. 2. The shaded regions above the curves depict the excluded part of the parameter space. The vertically shaded grey regions depict the excluded range of $m_{\chi}$ for the TA Amaterasu event as discussed in Sec. II.
@@
In Fig. 2, the constraints on $\tau_{\chi}^{-1}$ from TA’s UHECR data, TA’s gamma-ray flux upper limit and PAO’s gamma-ray flux upper limit are shown by the blue, magenta and red curves, respectively. As the bounds from non-observation of UHE neutrinos are orders of magnitude smaller than the bounds from UHECR and gamma-rays, the neutrino bounds are not shown in Fig. 2. In the top panel, the left plot shows the bounds on SHDM lifetime obtained from $\chi\to hh,~{}W^{+}W^{-},~{}ZZ$, the middle plot shows the bounds from $\chi\to b\bar{b},~{}t\bar{t},~{}c\bar{c}$, and the right plot shows the bounds from $\chi\to u\bar{u},~{}d\bar{d},~{}s\bar{s}$. Similarly, in the bottom panel, the left plot displays the bounds from the decay modes $\chi\to e^{+}e^{-},~{}\mu^{+}\mu^{-},~{}\tau^{+}\tau^{-}$, the middle plot from $\chi\to\nu_{e}\bar{\nu}_{e},~{}\nu_{\mu}\bar{\nu}_{\mu},~{}\nu_{\tau}\bar{\nu}_{\tau}$, and the right plot from $\chi\to\gamma\gamma,~{}gg$. The excluded regions of the $\tau_{\chi}^{-1}$ vs $m_{\chi}$ parameter in the plots correspond to flux larger than the experimental data or limits and hence disallowed.
@@
In all decay modes, for values of $m_{\chi}$ above $\sim 10^{13}$ GeV, the TA UHECR constraint curves cross the UHE gamma-ray constraint ones, opening a tiny parameter space supporting the SHDM origin of UHECRs. However, this inference is in conflict with the gamma-ray constraints from PAO except for the top left plot (bosonic decay modes) in Fig. 2. In this case, a small region above $m_{\chi}>10^{14}$ GeV is still allowed. However, the allowed range of $m_{\chi}$ for the TA Amaterasu event is found to be between $\sim(10^{11.8}-10^{13.6})$ GeV. Note the slight differences in the allowed mass ranges for different decay modes in the vertically shaded grey bands. Outside this range the SHDM decay can not explain both the TA Amaterasu event and the UHECR limit. Because, in this $m_{\chi}$ range, the constraints from the gamma-ray limits of TA and PAO disfavour the SHDM origin of TA Amaterasu event.
@@
The constraints extrapolated to higher $m_{\chi}$ values from the TA Amaterasu event allow a parameter space of $10^{-30}s^{-1}\lesssim\tau_{\chi}^{-1}\lesssim 10^{-29}~{}\rm s^{-1}$ for $m_{\chi}\gtrsim 10^{14}$ GeV, for SHDM originated UHECRs at the extreme energies of the spectrum. However, this UHE secondary flux being a diffuse flux, the PAO should also observe gamma-rays at similar energies. For instance, if we consider decay via the $b\bar{b}$ channel, the PAO can be expected to detect a UHE gamma-ray flux of about $(10^{-3}-10^{-4})$ $\rm km^{-2}~{}yr^{-1}~{}sr^{-1}$ at $\sim 10^{20}$ GeV. Hence, future detection of UHECRs at or above this energy will be crucial for constraining UHECRs from SHDM decay. Finally, we also study the dependence of these constraints on the choice of Galactic DM density profile and find that the results are nearly independent of density profile choice (see Appendix A).
@@ IV Discussion and Conclusion
To conclude, we have provided an updated estimate of multi-messenger constraints on the lifetime of SHDM by considering the recent TA Amaterasu event in conjunction with the TA UHECR data and non-observation of UHE gamma-rays and UHE neutrinos. While doing so, we have ensured consistency with the existing constraints in the literature [41, 62, 61, 59, 73, 40, 118]. Note that the constraints for heavier DM mass, i.e., $m_{\chi}>10^{12}$ GeV are not very robust due to the lack of data (UHECRs, UHE gamma-rays and UHE neutrinos) at energies above $\sim 10^{11}$ GeV. For harder UHECR flux anticipated from SHDM [41] or conventional diffusive acceleration [43, 82], the resulting constraints on the lifetime for $m_{\chi}\gtrsim 10^{12}$ GeV will be smaller than the ones obtained in this work. Nevertheless, such shorter lifetimes for the SHDM are not supported by UHE gamma-ray observations and do not impact the conclusion of this work.
@@ Acknowledgements.
This research (SC and PM) was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Understanding the Universe Through Neutrinos (code: ICTS/Neus2024/04). The work of D.B. is supported by the Science and Engineering Research Board (SERB), Government of India grants MTR/2022/000575 and CRG/2022/000603. D.B. also acknowledges the support from the Simons Foundation (Award Number:1023171-RC) to visit the International Institute of Physics, Natal, Brazil in May 2024 when part of this work was completed. SC and PS have received funding from SERB projects CRG/2021/002961 and MTR/2021/000540. PS thanks Toshihiro Fujii for providing crucial information on the TA observation and Pritam Das for useful discussion. The work of ND is supported by the Ministry of Education, Government of India via the Prime Minister’s Research Fellowship (PMRF) December 2021 scheme.
Appendix A Dependence on DM density profile
Figure A.1: Top plot: DM energy density vs Galactocentic distance for three different DM profiles. Bottom plot: Constraints on DM lifetime from UHECR vs DM mass for different DM profiles considering $b\bar{b}$ channel. The constraints are nearly independent of the DM profiles.

### v2

@@ III.1 Present UHE telescopes
In this sub-section, we present the constraints on SHDM lifetime from different observations of UHECRs, UHE gamma-rays and UHE neutrinos and analyse the origin of the TA Amaterasu event. The constraints are obtained by normalising the flux of these particles with the experimental data or limits from non-observation. For UHECRs, we consider the spectrum measured by TA together with the recent TA Amaterasu event as depicted in the left plot of Fig. 1. Including the UHECR data from other experiments such as PAO, AGASA and HiRes will yield similar values of DM lifetime as the CR spectral shapes from these experiments are also very similar to that of TA. For the UHE gamma-rays, we take the upper limits on the flux from TA [61, 62] and PAO [93, 57]. Whereas for the UHE neutrinos, we consider the upper limits on the flux from PAO [107] and IceCube [66]. We consider a wide range of $m_{\chi}=(10^{11}-10^{15})$ GeV and different two-body decay modes. For the decay modes, we take all the quarks (u, d, s, b, t, c), the gauge bosons (W, Z, $\gamma$, gluon), and the Higgs boson (h). The lower limit of $m_{\chi}$ is chosen to be $10^{11}$ GeV as our primary goal is to analyse the origin of the TA Amaterasu event at $2.44\times 10^{11}$ GeV. On the other hand, the upper limit of $m_{\chi}=10^{15}$ GeV is taken to be close to the GUT scale. For the given $m_{\chi}$, we constrain the lifetime such that the UHE secondary particle observations, limits and sensitivities are satisfied. In particular, we follow convention of exclusion plots and constrain the inverse of the lifetime, i.e., $\tau_{\chi}^{-1}$ rather than the lifetime ($\tau_{\chi}$). The constraints on $\tau_{\chi}^{-1}$ as a function of $m_{\chi}$ for different decay modes are depicted in Fig. 2. The shaded regions above the curves depict the excluded part of the parameter space. The vertically shaded gray regions depict the excluded range of $m_{\chi}$ for the TA Amaterasu event as discussed in Sec. II.
@@
In Fig. 2, the constraints on $\tau_{\chi}^{-1}$ from TA’s UHECR data, TA’s gamma-ray flux upper limit and PAO’s gamma-ray flux upper limit are shown by the blue, magenta and red curves, respectively. As the bounds from non-observation of UHE neutrinos are orders of magnitude smaller than the bounds from UHECRs and gamma-rays, the neutrino bounds are not shown in Fig. 2. In the top panel, the left plot shows the bounds on SHDM lifetime obtained from $\chi\to hh,~{}W^{+}W^{-},~{}ZZ$, the middle plot shows the bounds from $\chi\to b\bar{b},~{}t\bar{t},~{}c\bar{c}$, and the right plot shows the bounds from $\chi\to u\bar{u},~{}d\bar{d},~{}s\bar{s}$. Similarly, in the bottom panel, the left plot displays the bounds from the decay modes $\chi\to e^{+}e^{-},~{}\mu^{+}\mu^{-},~{}\tau^{+}\tau^{-}$, the middle plot from $\chi\to\nu_{e}\bar{\nu}_{e},~{}\nu_{\mu}\bar{\nu}_{\mu},~{}\nu_{\tau}\bar{\nu}_{\tau}$, and the right plot from $\chi\to\gamma\gamma,~{}gg$. The excluded regions of the $\tau_{\chi}^{-1}$ vs $m_{\chi}$ parameter in the plots correspond to flux larger than the experimental data or limits and hence disallowed.
@@
In all decay modes, for values of $m_{\chi}$ above $\sim 10^{13}$ GeV, the TA UHECR constraint curves cross the UHE gamma-ray constraint ones, opening a tiny parameter space supporting the SHDM origin of UHECRs. However, this inference is in conflict with the gamma-ray constraints from PAO except for the top left plot (bosonic decay modes) in Fig. 2. In this case, a small region above $m_{\chi}>10^{14}$ GeV is still allowed for UHECRs of SHDM origin. However, the allowed range of $m_{\chi}$ for the TA Amaterasu event is found to be between $\sim(10^{11.8}-10^{13.6})$ GeV. Note the slight differences in the allowed mass ranges for different decay modes in the vertically shaded grey bands. The gamma-ray constraints being the strongest in this $m_{\chi}$ range, the SHDM origin of the TA Amaterasu event is found to be in significant conflict with the gamma-ray constraints. Whereas, outside this range the SHDM decay can not explain both the TA Amaterasu event and the UHECR constant flux limit. Because, in this $m_{\chi}$ range, the constraints from the gamma-ray limits of TA and PAO disfavour the SHDM origin of TA Amaterasu event.
@@
The constraints extrapolated to higher $m_{\chi}$ values from the TA Amaterasu event allow a parameter space of $10^{-30}s^{-1}\lesssim\tau_{\chi}^{-1}\lesssim 10^{-29}~{}\rm s^{-1}$ for $m_{\chi}\gtrsim 10^{14}$ GeV, for a SHDM originated UHECR flux at the extreme energies of the spectrum. However, this UHE secondary flux being a diffuse flux, the PAO should also observe gamma-rays at similar energies. For instance, if we consider decay via the $b\bar{b}$ channel, the PAO can be expected to detect a UHE gamma-ray flux of about $(10^{-3}-10^{-4})$ $\rm km^{-2}~{}yr^{-1}~{}sr^{-1}$ at $\sim 10^{20}$ GeV. Hence, future detection of UHECRs at or above this energy will be crucial for constraining UHECRs from SHDM decay. We also study the dependence of these constraints on the choice of Galactic DM density profile and find that the results are nearly independent of density profile choices (see Appendix A). Large DM density at the Galactic center can yield additional constraints on the Amaterasu’s SHDM origin. However, these constraints are weak in comparison to those from the diffuse UHECR flux due to certain uncertainties which are discussed in detail in Appendix B.
@@ IV Discussion and Conclusion
To conclude, we have provided an updated estimate of multi-messenger constraints on the lifetime of SHDM by considering the recent TA Amaterasu event in conjunction with the TA UHECR data and non-observation of UHE gamma-rays and UHE neutrinos. While doing so, we have ensured consistency with the existing constraints in the literature [45, 70, 69, 67, 81, 44, 127, 85]. Note that the constraints for heavier DM mass, i.e., $m_{\chi}>10^{12}$ GeV are not very robust due to the lack of data (UHECRs, UHE gamma-rays and UHE neutrinos) at energies above $\sim 10^{11}$ GeV. For harder UHECR flux anticipated from SHDM [45] or conventional diffusive acceleration [51, 91], the resulting constraints on the lifetime for $m_{\chi}\gtrsim 10^{12}$ GeV will be smaller than the ones obtained in this work. Nevertheless, such shorter lifetimes for the SHDM are not supported by UHE gamma-ray observations and do not impact the conclusion of this work.
@@ Acknowledgements.
PS thanks Toshihiro Fujii for providing crucial information on the TA observation and Pritam Das, Ranjan Laha for useful comments. We thank Subir Sarkar for valuable comments and for pointing out important references related to the work. This research (SC and PM) was supported in part by the International Centre for Theoretical Sciences (ICTS) for participating in the program - Understanding the Universe Through Neutrinos (code: ICTS/Neus2024/04). The work of ND is supported by the Ministry of Education, Government of India via the Prime Minister’s Research Fellowship (PMRF) December 2021 scheme. The work of DB is supported by the Science and Engineering Research Board (SERB), Government of India grants MTR/2022/000575 and CRG/2022/000603. DB also acknowledges the support from the Simons Foundation (Award Number:1023171-RC) to visit the International Institute of Physics, Natal, Brazil in May 2024 when part of this work was completed. SC and PS have received funding from SERB projects CRG/2021/002961 and MTR/2021/000540. PM acknowledges the use of HPC cluster at SPS, JNU funded by DST-FIST.
Appendix A Dependence on DM density profile on diffuse flux
Figure A.1: Left plot: DM energy density vs Galactocentic distance for three different DM profiles. Right plot: Constraints on DM lifetime from diffuse UHECR flux vs DM mass for different DM profiles considering $b\bar{b}$ channel. The constraints are nearly independent of the DM profiles. The region above the lines are excluded.
@@ Appendix A Dependence on DM density profile on diffuse flux
Appendix B Galactic Center contribution
Figure B.1: Left plot : constraints on SHDM lifetime from the direction of Galactic center for three different DM profiles : NFW (solid blue), Burkert (dotted red) and Einesto (dashed green). Right plot : Variations on constraints due to attenuation of UHECR from the GC shown by solid blue (without attenuation) and thick dot-dot-dashed blue (attenuated with $n_{\rm CMZ}\sim 10^{4}$ cm-3) lines. For comparison, diffuse flux bound is plotted with dot-dashed blue line. The DM profile considered here is the NFW profile.
Figure B.2: Constraints on SHDM lifetime from the direction of Amaterasu (orange dot-dot dashed) compared with that from the diffuse flux (blue solid) for NFW DM profile.
In the above calculations, we focus on the diffuse part of the secondary fluxes originating from SHDM decay in the Milky way. However, significant anisotropy [129] in these secondary fluxes is expected due to our location away from the Galactic center (GC). The anisotropy is expected due to the large DM density in the Galactic center. Indeed, this DM density at the GC can vary for different DM profiles. To demonstrate this, we again consider the Einesto [88] and the Burkert [90] models, together with the NFW profile. To estimate this contribution from the Galactic center direction, we consider a solid angle defined by the Galactic latitude ($-5^{\circ},5^{\circ}$) and longitude ($0^{\circ},5^{\circ}$). The resulting constraints on the SHDM lifetime for the three DM profiles are shown in the left panel of Fig. B.1. The Einesto’s (dashed green) and Burkert’s (dotted red) DM profiles yield the strongest and weakest constraints, respectively. The constraint from the NFW profile (solid blue) remains slightly weaker than those from the Einesto profile. Clearly, in comparison to the DM model dependence of the diffuse flux (Fig. A.1) the GC flux shows a large uncertainty and hence not the preferred choice for phenomenology.
In addition to the DM model uncertainties, the flux from the GC region might be affected by the associated attenuation of UHECRs due to the large density of molecular cloud at the Milky Way center. The flux from the Galactic center would suffer attenuation due to the interaction (e.g., inelastic $pp$ collision) with the dense molecular gas in the “Central Molecular Zone (CMZ)”. Depending on the density, the CMZ extending up to the radius of about $R=300$ pc [130] can cause substantial attenuation to this secondary fluxes. In particular, the CMZ density ($n_{\rm CMZ}$) can vary in several orders, the difference arises in different length scales around the Galactic center. Typically, the $n_{\rm CMZ}$ is quoted in the range $\sim(10^{2}-10^{3})~{}\rm cm^{-3}$. However, there are regions of dense molecular gas clouds within the CMZ with density exceeding $10^{3}~{}\rm cm^{-3}$ [130, 132]. Indeed, in the parsec scale the density of molecular gas clouds in the CMZ can reach up to (${}\sim 10^{4}-10^{7}~{}\rm cm^{-3}$) [131]. To mimic the absorption effect of the high density gas clouds, we extend the upper limit of CMZ density to $10^{4}~{}\rm cm^{-3}$. For these high densities the absorption of UHECRs from GC is found to be non negligible and may lead to larger attenuation than that of the typical CMZ density. Thus, along with the DM model, this uncertainty of the scale and density of the $n_{\rm CMZ}$ adds to the ambiguity of the SHDM flux from GC.
This has been shown in the plot on the right of Fig. B.1 shows the limit on $\tau^{-1}_{\chi}$ from the Galactic center region for NFW profile with decay mode $\chi\to b\bar{b}$. The bound without the CMZ attenuation is shown by the continuous solid blue curve. The bound with attenuation corresponding to $n_{\rm CMZ}=10^{4}~{}\rm cm^{-3}$ is shown by the thick dot-dot dashed blue curves. We also show the diffuse flux bound by the dashed-dot blue curve for comparison. This shows that the attenuation of UHECRs from the GC can be significant for the large CMZ density $n_{\rm CMZ}\sim 10^{4}~{}\rm cm^{-3}$. The attenuation is found to be negligible for $n_{\rm CMZ}\leq 10^{3}~{}\rm cm^{-3}$. The attenuation can also be impacted by the composition of the CMZ. For our estimation, we choose one nucleon per molecule to be conservative. Note, that the UHECR protons can also interact with low energy photons (photo-pion production), but the attenuation due to this process is found to be negligible compared to the $pp$ one.
The uncertainty in the UHECR flux from the GC is dominated by the uncertainty due to the DM profile, but may also have contribution from the CMZ. Therefore, the GC flux is not the best candidate for such phenomenological studies. Hence, we focus on diffuse contribution of the SHDM to obtain the bounds on the decay rate. Keeping these model uncertainties aside, the non-discovery of these fluxes in PAO should still be important.
Finally, we also compute the constraints on SHDM lifetime from the direction of the Amaterasu by varying the Galactic coordinates in the uncertainty range of the Amaterasu’s arrival direction for the NFW profile. Note that the constraints are almost independent of the choice of DM profile as the Amaterasu’s arrival direction is about $45^{\circ}$ away from the Galactic center direction. The resulting constraints are compared with those from diffuse flux and the Galactic center direction (NFW) in Fig. B.2. The constraints from the Amaterasu’s direction (orange dot-dot dashed) are slightly stronger (about a factor of $1.5$) than those from the diffuse flux. Therefore, our constraints derived from the diffuse UHECR flux are the conservative ones.